OpenStudy (anonymous):

at a certain baseball diamond the distance form home plate through the pitcher rubber to the far well (to dead center) is 361 ft given that the infield forms a square 90 ft on each side how far is it form third base to dead center round your answer to nearest tenth

OpenStudy (ranga):



OpenStudy (ranga):

H is Home Plate, A is First Base. T is Third Base. C is Dead Center. HA = HT = 90 ft. ; HC = 361 ft. Angle THA = 90 degrees. So angle THC = 45 degrees What is TC? In triangle TCH use the law of cosines to find TC. TC^2 = HT^2 + HC^2 - 2(HT)(HC)cos(THC) = 90^2 + 361^2 - 2(90)(361)cos(45) Find TC.

OpenStudy (anonymous):

how i do that sir

OpenStudy (anonymous):

sir u theie

OpenStudy (anonymous):

i got 304

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

sir jii where r uhelp plz

OpenStudy (ranga):

I am getting 304 ft too.

OpenStudy (anonymous):

so is that it

OpenStudy (anonymous):

or do i need to find more angle

OpenStudy (anonymous):

is that is or do i need to find mmore

OpenStudy (ranga):

I see only one question: "how far is it form third base to dead center round your answer to nearest tenth" They want the answer to the nearest tenth: 304.1 feet.

OpenStudy (anonymous):

ok thanks you sir